# Physical meaning of non-extensive term in Massieu functions

### v2

@@ Abstract
In this paper, we explore the significance of non-extensive terms in Massieu functions. We demonstrate that these terms, of order $\ln N$, where $N$ represents the number of elements, can be expressed in relation to thermodynamic fluctuations. We focus on systems large enough to mitigate finite size effects, concentrating on the $O(\ln N)$ difference present in both ideal and non-ideal scenarios. The analytical verification in ideal cases and numerical validation in non-ideal scenarios corroborate our results. The examples used are hard sphere and hard disk fluids and a one-dimensional spin lattice, emphasizing the applicability of the results across different classes of systems.
@@ I Introduction
Non-extensive terms in the Boltzmann’s definition of entropy are usually neglected in the thermodynamic limit, where the results of statistical mechanics converge to the ones of thermodynamics. The differences between thermodynamic and statistical mechanics extensive quantities are of order $\ln N$, where $N$ is the number of elements that compose the system. The logarithmic relationship stems from the number of possible permutations of $N$ particles: $N!$. When we take the logarithm of $N!$, it includes a term proportional to $\ln N$. This term can become significant, especially in molecular simulations of finite systems employing periodic boundary conditions Landau and Binder (2005).
@@
We present a general approach to demonstrate that those differences can be written in terms of thermodynamic fluctuations, that, in turn, can be obtained from thermodynamic quantities. This means that the thermodynamic equation of state contains the information needed to calculate non-extensive terms of statistical mechanics quantities. Or, the other way around, non-extensive terms of statistical mechanics quantities have information about thermodynamic fluctuations. Our analysis is restricted to systems large enough so that finite size effects, analyzed in, for example, Hill (1964); Bedeaux et al. (2020), other than the logarithmic term, can be neglected.
@@
The difference $O(\ln N)$ is present in the non-ideal and ideal (negligible interactions) cases. Excess extensive quantities are obtained when the ideal value is subtracted. It can be seen that thermodynamic and statistical mechanics excess quantities differ by a term that is $O(N^{0})$. For example, the excess free energy of a system of particles, calculated using the partition function, includes a non-extensive term $O(N^{0})$; and the excess chemical potential has a non-extensive term $O(N^{-1})$. This term was calculated using the Widom insertion formula in Ref. Siepmann et al. (1992), where the result is presented as a finite size correction to numerical calculation of the excess chemical potential in a system with periodic boundary conditions. An equivalent expression was derived with a different procedure in Ref. Di Muro and Hoyuelos (2021), where it was shown that the non-extensive term is relevant for the evaluation of transition rates and, in turn, of the diffusion coefficient. Actually, the relationship between self-diffusion and collective diffusion coefficients, known as one of the the Darken equations Darken (1948), was derived in Di Muro and Hoyuelos (2021) using the non-extensive term of the excess chemical potential. Here we present a general approach, valid for any Massieu function, that includes the mentioned results; that is, it is not restricted to the analysis of the free energy. It is generally more convenient to use Massieu functions instead of thermodynamic potentials in order to analyze fluctuations (see (Callen, 1985, Ch. 19)); there are simple relationships that connect them.
@@
In the next sections we present the basic equations and notation (Sec. II), the calculation of the difference between thermodynamic and statistical mechanics quantities (Sec. III), applications of the results to a system of particles and a spin lattice (Sec. IV and V), and the conclusions (Sec. VI).
@@ II Basic equations and notation
Let us consider a system described by the extensive variables $\vec{X}=(X_{1},X_{2},\dots)$. The typical example is $\vec{X}=(U,V,N)$, where $U$, $V$ and $N$ are the internal energy, volume and particle number. The Boltzmann entropy is
@@
First, we present the definition of Massieu functions in their thermodynamic (TH) and statistical mechanics (SM) versions for a specific situation: a particle sub-system of volume $V$ that is part of a larger complex; the complement of the sub-system plays the role of a reservoir that imposes a temperature $T$ and a chemical potential $\mu$. Local fluctuations in $U$ and $N$ are present, and $V$ is a constant parameter by definition. We consider that the vector $\vec{X}=(U,N)$ contains the extensive variables, but there can be other parameters, like $V$ in this case, that are fixed by external constraints. TH Massieu functions are obtained from successive applications of Legendre transforms, starting from the entropy. For example:
@@
Now we turn to a statistical mechanics description. We have to specify how to calculate Massieu functions in this context. Instead of the Legendre transform, the mathematical transformation that generates SM Massieu functions starting from the entropy is the Laplace transform. Quantities with tilde, in the statistical mechanics context, contain non-extensive terms; we are interested in, for example, the difference $\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-S[\tfrac{1}{T},\tfrac{\mu}{T}]$ at order $\ln N$.
The TH Massieu functions $S[\tfrac{1}{T},N]$ and $S[\tfrac{1}{T},\tfrac{\mu}{T}]$, equal to $-F/T$ and $PV/T$, correspond to the logarithm of the canonical and grand canonical partition functions, $\mathcal{Z}(\tfrac{1}{T},N)$ and $\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})$, respectively, as long as non-extensive terms are neglected in the thermodynamic limit. The SM Massieu functions are given by
@@
$P(\omega)=\frac{\exp[-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}]}{\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})}=\exp[-\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}].$ (9)
@@
Taking the sum of $P(\hat{U},\hat{N})$ over possible values of $\hat{U}$ we obtain the probability of $\hat{N}$:
@@
In general, for a given microstate $\omega$ the extensive variables take the values $\hat{\vec{X}}=(\hat{X}_{1},\dots,\hat{X}_{r})$; the list does not include extensive parameters fixed by external constraints; variables without hat, $X_{i}$, are the thermodynamic averages. The probability of the microstate $\omega$ is given by (see, for example, (Callen, 1985, Ch. 19))
@@
If we have $s$ extensive variables and $r-s$ intensive parameters, that we can write as $\vec{X}^{s}=(X_{1},\cdots,X_{s})$ and $\vec{Y}^{r-s}=(Y_{s+1},\cdots,Y_{r})$, then the SM Massieu function is the logarithm of the partition function in the corresponding ensemble:
@@
where we have used the average values of the extensive variables (without hat), but the same expression holds for $\hat{\vec{X}}^{s}$ for a given microstate. Also, the SM Massieu function that depends on $s$ extensive variables and $r-s$ intensive parameters is given by $r-s$ Laplace transforms of $\tilde{S}(\hat{\vec{X}})$:
@@
$\sum_{\hat{X}_{s+1}}\cdots\sum_{\hat{X}_{r}}$ $e^{\tilde{S}(\hat{\vec{X}})}e^{-Y_{s+1}\hat{X}_{s+1}\dots-Y_{r}\hat{X}_{r}}.$
@@
III Non-extensive term and thermodynamic fluctuations
@@ III.1 $\hat{U},\hat{N},V$ system
Terms $D(N_{\rm max})$ and $D(0)$ were canceled since the probability distribution vanishes at the extreme values of the variable.
Then, from (17) and (III.1) we have that
@@
The average can be written in terms of mean squared fluctuations if we expand $\frac{\partial\tilde{S}[\frac{1}{T},\hat{N}]}{\partial\hat{N}}$ in a Taylor series around the mean value $N$:
@@
where $\Delta\hat{N}=\hat{N}-N$, and terms $O(N^{-2})$ are neglected [the mean squared or higher order fluctuations are $O(N)$, and the $i$-th derivative is $O(N^{-i+1})$]; the third derivative of $\tilde{S}$ was replaced by the third derivative of $S$ (without tilde) since the difference contributes with a term $O(N^{-2})$. Combining the last two equations,
$\frac{\partial}{\partial N}[\tilde{S}[\tfrac{1}{T},N]-S[\tfrac{1}{T},N]]=-\frac{1}{2}\frac{\partial^{3}S[\tfrac{1}{T},N]}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle.$ (23)
@@
where $c$ is an integration constant ($c$ may be neglected respect to the logarithm of fluctuations $O(\ln N)$).
@@
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}+c,$ (28)
The excess Massieu function (TH or SM) is defined as the difference $S_{\text{ex}}=S-S_{\text{id}}$ (with or without tilde). Doing subtraction between Eqs. (27) and (28), we obtain,
@@
Then, the non-extensive term of the excess part does not behave as $\ln N$ but, instead, it is $O(N^{0})$.
@@ III.2 General derivations
The average of its derivative respect to $\hat{X}_{i}$, with $i=1,\cdots,s$, is
@@ A Taylor expansion around $\vec{X}^{s}$ is used to evaluate the average:
$\left\langle\frac{\partial\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}{\partial\hat{X}_{i}}\right\rangle$ $=\frac{\partial\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}}+$
$\frac{1}{2}\sum_{j,k}\frac{\partial^{3}S[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle,$ (33)
@@
Let us call $C$ the (symmetric) fluctuation matrix. Component $C_{jk}$ is given by (see Callen (1985))
@@
$=\frac{1}{2}{\rm tr}\left(\frac{\partial C^{-1}}{\partial X_{i}}\cdot C\right),\qquad i=1,\cdots,s$ (36)
@@
Subtracting the last two equations, we obtain the non-extensive term of the excess Massieu function:
@@
Using the result of Eq. (37), and the normalization of the probability distribution of fluctuations, it can be shown that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant; see Appendix A. It is known that this constant is zero for some solvable systems, meaning that $\tilde{S}[\vec{Y}]$ is exactly extensive; this is the case for the one-dimensional spin lattice, as is mentioned in Sec. V.2 below, and for the grand canonical ensemble of an ideal gas (see Sec. 8.2 in Swendsen (2017), where the Massieu function that depends on $\vec{Y}=(\frac{1}{T},\frac{\mu}{T})$ is called the grand canonical entropy).

### v3

# Physical meaning of nonextensive term in Massieu functions

@@ Abstract
In this paper we explore the significance of nonextensive terms in Massieu functions. Finite-size effects are in many cases dominated by a term proportional to the surface area. Nevertheless, in numerical simulations of finite systems with periodic boundary conditions, the nonextensive term can become the dominant correction to Massieu functions. This paper presents a general approach linking these nonextensive terms to thermodynamic fluctuations, demonstrating that equations of state inherently encode this information. Numerical simulations corroborate our results. The examples used are hard sphere and hard disk fluids and a one-dimensional spin lattice, emphasizing the applicability of the results across different classes of systems.
@@ I Introduction
Nonextensive terms in the Boltzmann definition of entropy are usually neglected in the thermodynamic limit, where the results of statistical mechanics converge to those of thermodynamics. The differences between extensive quantities of thermodynamic and statistical mechanics are of the order $\ln N$, where $N$ is the number of elements that compose the system. The logarithmic relationship stems from the number of possible permutations of $N$ elements: $N!$. When we take the logarithm of $N!$, it includes a term proportional to $\ln N$. This term can become significant, especially in molecular simulations of finite systems employing periodic boundary conditions Landau and Binder (2005); in the examples of small systems analyzed here, this term has a contribution between 4% and 10% for excess quantities (relative to the ideal value).
@@
We present a general approach based on Massieu functions to demonstrate that these differences can be written in terms of thermodynamic fluctuations that, in turn, can be obtained from thermodynamic quantities. This means that the thermodynamic equation of state contains the information needed to calculate nonextensive terms of statistical mechanics quantities. Or, the other way around, nonextensive terms of statistical mechanics quantities have information about thermodynamic fluctuations.
Let us call $X_{\text{finite}}$ a Massieu function or a thermodynamic potential of a finite system and $X$ its value at the thermodynamic limit; $X_{\text{finite}}$ includes other corrections in addition to the logarithmic one. The most important is proportional to the surface area of the system, $L^{2}\sim V^{2/3}\sim N^{2/3}$, where $L$ is the system size, $V$ is the system volume, and $N$ is the particle number; see, for example, Sec. 5.1 in Bedeaux et al. (2020). Then,
$X_{\text{finite}}-X=O(\ln N)+O(N^{2/3}).$ (1)
(Additional possible corrections, of order $N^{1/3}$ and $N^{0}$, that correspond to the effects of edges and corners, are not considered here.) In order to determine which of the two corrections in (1) is dominant, we have to compare the system size with the correlation length of fluctuations Lebowitz and Percus (1961). For a real system with strong interactions that produce a long correlation length similar to the system size, the correction $O(\ln N)$ can be neglected with respect to surface effects; this is the situation studied in the theory introduced by Hill Hill (1964); Bedeaux et al. (2020). Periodic boundary conditions are often used to simulate bulk properties, but despite the absence of physical surfaces, finite-size effects can still contribute corrections dominant respect to the term $O(\ln N)$; they are usually called implicit or anomalous finite-size effects Román et al. (2008). Salacuse et alSalacuse et al. (1996a) demonstrated that if the pair distribution function decays rapidly with distance, then implicit finite-size corrections can be neglected, a criterion equivalent to considering the absence of long-range correlations Lebowitz and Percus (1961); in this case the logarithmic correction is dominant. The extreme situation is the ideal system, where interactions can be neglected and the correlation length is zero; for the ideal case with periodic boundary conditions, the only correction is $O(\ln N)$.
Our analysis is restricted to systems with a correlation length smaller than $L$, so implicit finite-size effects can be neglected. The corrections obtained in this case are called explicit finite-size effects Román et al. (2008). For example, we can analyze the free energy of a dilute gas, with a fixed number of particles $N$, and progressively increase the density by reducing the volume $V$, with periodic boundary conditions. Initially, the system is ideal, interactions can be neglected, and our assumption for the correlation length holds. As the density is increased, interactions become more relevant and the correlation length also increases. A point will be reached for which the correlation length is similar to $L$ and implicit finite-size effects cannot be neglected anymore. At this point, our present analysis does not hold, but there is a range of concentration values, starting from the dilute gas to this point, where our results correctly predict the correction $O(\ln N)$. Actually, for nonideal systems, it is more convenient to analyze excess quantities, that is, for example, the free energy minus its ideal value. It can be seen that, for excess quantities, the correction is $O(N^{0})$ instead of $O(\ln N)$.
@@
The excess free energy of a system of particles, calculated using the partition function, includes the nonextensive term mentioned $O(N^{0})$, and the excess chemical potential has a correction term $O(N^{-1})$. This term was calculated using the Widom insertion formula in Ref. Siepmann et al. (1992), where the result is presented as a finite-size correction to numerical calculation of the excess chemical potential in a system with periodic boundary conditions. An equivalent expression was derived with a different procedure in Ref. Di Muro and Hoyuelos (2021), where the nonextensive term was shown to be relevant for the evaluation of transition rates and, in turn, of the diffusion coefficient. In fact, the relationship between self-diffusion and collective diffusion coefficients, known as one of the Darken equations Darken (1948), was derived in Di Muro and Hoyuelos (2021) using the correction term for the excess chemical potential. Here we present a general approach, valid for any Massieu function, that includes the mentioned results; that is, it is not restricted to the analysis of the free energy. It is generally more convenient to use Massieu functions instead of thermodynamic potentials to analyze fluctuations (see (Callen, 1985, Ch. 19)); there are simple relationships that connect them.
Therefore, we study corrections order $\ln N$ in thermodynamic quantities (Massieu functions or thermodynamic potentials); the procedure cannot be applied to calculate corrections to statistical mechanics quantities involving microscopic information, such as the radial distribution function. Corrections to the structure factor or the radial distribution function of a system of particles were analyzed in Salacuse et alSalacuse et al. (1996a), based on the theory developed by Lebowitz and Percus Lebowitz and Percus (1961); it has been applied, for example, to evaluate the structure factor for a model krypton fluid Salacuse et al. (1996b) or fluctuations in the number of hard disks and hard parallel squares deposited on a finite surface Román et al. (1997). The main feature of the present approach is the possibility to obtain a general result for the nonextensive term of any Massieu function without using the radial distribution function. We demonstrate that the nonextensive term is directly related to the determinant of the equilibrium fluctuation matrix; see Eq. (40) below. We illustrate the generality of the approach applying the results to particle systems and to a spin lattice.
@@
In the next sections we present the basic equations and notation (Sec. II), the calculation of the difference between thermodynamic and statistical mechanics quantities (Sec. III), applications of the results to a system of particles and a spin lattice (Secs. IV and V), and the conclusions (Sec. VI).
@@ II Basic equations and notation
Let us consider a system described by the extensive variables $\vec{X}=(X_{1},X_{2},\dots)$. The typical example is $\vec{X}=(U,V,N)$, where $U$, $V$, and $N$ are the internal energy, volume, and particle number. The Boltzmann entropy is
@@
First, we present the definition of Massieu functions in their thermodynamic (TH) and statistical mechanics (SM) versions for a specific situation: a particle subsystem of volume $V$ that is part of a larger complex; and the complement of the subsystem plays the role of a reservoir that imposes a temperature $T$ and a chemical potential $\mu$. Local fluctuations in $U$ and $N$ are present, and $V$ is a constant parameter by definition. We consider that the vector $\vec{X}=(U,N)$ contains the extensive variables, but there can be other parameters, like $V$ in this case, that are fixed by external constraints. TH Massieu functions are obtained from successive applications of Legendre transforms, starting from the entropy. For example,
@@
Now we turn to a statistical mechanics description. We have to specify how to calculate Massieu functions in this context. Instead of the Legendre transform, the mathematical transformation that generates SM Massieu functions starting from the entropy is the Laplace transform. Quantities with tilde, in the statistical mechanics context, contain nonextensive terms; we are interested in, for example, the difference $\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-S[\tfrac{1}{T},\tfrac{\mu}{T}]$ at order $\ln N$.
The TH Massieu functions $S[\tfrac{1}{T},N]$ and $S[\tfrac{1}{T},\tfrac{\mu}{T}]$, equal to $-F/T$ and $PV/T$, correspond to the logarithm of the canonical and grand canonical partition functions, $\mathcal{Z}(\tfrac{1}{T},N)$ and $\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})$, respectively, as long as nonextensive terms are neglected in the thermodynamic limit. The SM Massieu functions are given by
@@
$P(\omega)=\frac{\exp[-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}]}{\mathcal{Z}(\tfrac{1}{T},\tfrac{\mu}{T})}=\exp(-\tilde{S}[\tfrac{1}{T},\tfrac{\mu}{T}]-\tfrac{1}{T}\hat{U}+\tfrac{\mu}{T}\hat{N}).$ (10)
@@
Taking the sum of $P(\hat{U},\hat{N})$ over possible values of $\hat{U}$, we obtain the probability of $\hat{N}$:
@@
In general, for a given microstate $\omega$ the extensive variables take the values $\hat{\vec{X}}=(\hat{X}_{1},\dots,\hat{X}_{r})$; the list does not include extensive parameters fixed by external constraints; variables without a hat, $X_{i}$, are the thermodynamic averages. The probability of the microstate $\omega$ is given by (see, for example, (Callen, 1985, Ch. 19))
@@
If we have $s$ extensive variables and $r-s$ intensive parameters, that we can write as $\vec{X}^{s}=(X_{1},\dots,X_{s})$ and $\vec{Y}^{r-s}=(Y_{s+1},\dots,Y_{r})$, then the SM Massieu function is the logarithm of the partition function in the corresponding ensemble:
@@
where we have used the average values of the extensive variables (without a hat), but the same expression holds for $\hat{\vec{X}}^{s}$ for a given microstate. Also, the SM Massieu function that depends on $s$ extensive variables and $r-s$ intensive parameters is given by $r-s$ Laplace transforms of $\tilde{S}(\hat{\vec{X}})$:
@@
$\sum_{\hat{X}_{s+1}}\dots\sum_{\hat{X}_{r}}$ $e^{\tilde{S}(\hat{\vec{X}})}e^{-Y_{s+1}\hat{X}_{s+1}\dots-Y_{r}\hat{X}_{r}}.$
@@
III nonextensive term and thermodynamic fluctuations
@@ III.1 $\hat{U},\hat{N},V$ system
Terms $D(N_{\rm max})$ and $D(0)$ were canceled since the probability distribution vanishes at the extreme values of the variable. Then, from (18) and (III.1) we have that
@@
The average can be written in terms of mean-squared fluctuations if we expand $\frac{\partial\tilde{S}[\frac{1}{T},\hat{N}]}{\partial\hat{N}}$ in a Taylor series around the mean value $N$:
@@
where $\Delta\hat{N}=\hat{N}-N$, and the terms $O(N^{-2})$ are neglected [the mean squared or higher-order fluctuations are $O(N)$, and the $i$th derivative is $O(N^{-i+1})$]; the third derivative of $\tilde{S}$ was replaced by the third derivative of $S$ (without tilde) since the difference contributes with a term $O(N^{-2})$. Combining the last two equations,
$\frac{\partial}{\partial N}(\tilde{S}[\tfrac{1}{T},N]-S[\tfrac{1}{T},N])=-\frac{1}{2}\frac{\partial^{3}S[\tfrac{1}{T},N]}{\partial N^{3}}\langle(\Delta\hat{N})^{2}\rangle.$ (24)
@@
where $c$ is an integration constant [$c$ may be neglected respect to the logarithm of fluctuations $O(\ln N)$]. Let us notice that this expression was obtained by Lebowitz and Percus in terms of the free energy for a fluid particle system; see Eq. (3.11) in Ref. Lebowitz and Percus (1961). The present development based on Massieu functions allows the derivation of more general expressions, as shown in the next section.
@@
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln\langle(\Delta\hat{N})^{2}\rangle_{\text{id}}+c.$ (29)
The excess Massieu function (TH or SM) is defined as the difference $S_{\text{ex}}=S-S_{\text{id}}$ (with or without a tilde). Doing subtraction between Eqs. (28) and (29), we obtain,
@@
Then, the nonextensive term of the excess part does not behave as $\ln N$ but, instead, it is $O(N^{0})$.
@@ III.2 General derivations
The average of its derivative respect to $\hat{X}_{i}$, with $i=1,\dots,s$, is
@@ A Taylor expansion around $\vec{X}^{s}$ is used to evaluate the average:
$\left\langle\frac{\partial\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}{\partial\hat{X}_{i}}\right\rangle$ $=\frac{\partial\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}}$
$+\frac{1}{2}\sum_{j,k}\frac{\partial^{3}S[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle,$ (34)
@@
Let us call $C$ the (symmetric) fluctuation matrix. The component $C_{jk}$ is given by (see Callen (1985))
@@
$=\frac{1}{2}{\rm tr}\left(\frac{\partial C^{-1}}{\partial X_{i}}\cdot C\right),\qquad i=1,\dots,s$ (37)
@@
Subtracting the last two equations, we obtain the nonextensive term of the excess Massieu function:
@@
Using the result of Eq. (38), and the normalization of the probability distribution of fluctuations, it can be shown that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant; see Appendix A. It is known that this constant is zero for some solvable systems, meaning that $\tilde{S}[\vec{Y}]$ is exactly extensive; this is the case for the one-dimensional spin lattice, as is mentioned in Sec. V.2 below, and for the grand canonical ensemble of an ideal gas (See Sec. 8.2 in Swendsen (2017), where the Massieu function that depends on $\vec{Y}=(\frac{1}{T},\frac{\mu}{T})$ is called the grand canonical entropy.)